OpenStudy (surryyy):

simplify the expression: 3[9-(2+3)]^2

OpenStudy (anonymous):

first deal with the bracket

OpenStudy (anonymous):

2+3=5

OpenStudy (anonymous):

9-5=4

OpenStudy (anonymous):

now deal with the power first before the multipliation

OpenStudy (anonymous):

4^2=16

OpenStudy (anonymous):

16*3=48 answer

OpenStudy (surryyy):

@imqwerty

imqwerty (imqwerty):

3[9-(2+3)]^2 3[9-5]^2 3[4]^2 3(4x4) 3(16) =48